ID |
Date |
Author |
Type |
Category |
Subject |
4991
|
Tue Jul 19 20:36:08 2011 |
rana | Update | Computers | VirtualBox + Windows 7 on rossa |
I installed Virtual Box on rossa. Then I put Windows 7 in there and am now installing Altium.
You can run Windows on rossa by just clicking Applications -> System Tools -> Virtual Box. |
4990
|
Tue Jul 19 13:45:22 2011 |
Suresh | Update | SAFETY | Painful Voltage Regulator and Circuit-Testing Lessons |
There is should be a few IDC connectors in the lab (and some ribbon cable) using which you can proceed with the testing of the circuit, if you prefer. If not we can get them from our ever helpful electronics division at Downs. In any case there is no need to lose time waiting for parts to arrive.
|
4989
|
Tue Jul 19 10:54:14 2011 |
steve | Update | General | large sensor card can not be found |
Please return sensor card to laser log box so others can use it. We have only one larger fluorescent sensor card. |
4988
|
Tue Jul 19 10:18:24 2011 |
rana | Update | LSC | Big ol' mess |
Remember, as per our marker board conversation, that the resistor noise as excitation method only works if the gain of all of the channels is set to something high (like 45 dB).
At 0 dB, the resistor noise is only 30 nV/rHz, whereas the ADC noise is more like 10000 nV/rHz... |
4987
|
Tue Jul 19 09:19:14 2011 |
steve | Update | SUS | PRM damping restored |
The PRM watchdogs were tripped. The side was kicked up to 180mV Damping was restored. |
4986
|
Mon Jul 18 22:26:43 2011 |
Nicole | Update | SAFETY | Painful Voltage Regulator and Circuit-Testing Lessons |
Today I learned some important circuit-building lessons while testing my photosensor circuit box (i.e. how NOT to test a circuit and, conversely, things that should be done instead).
I blew my first circuit today. The victim is in the photo below (bottom 7805 voltage regulator). The plastic covering fell off after I removed the fried regulator. After checking various components, I figured out that I blew the circuit because I had forgotten to ground the regulator. Although this was very unfortunate, I did make an important discovery. While testing the voltage output of the 7805 voltage regulator (I put a new one), I discovered that contrary to the claims of the datasheet, an input voltage of 5V will not produce a steady 5V supply. I found that at 5V, my regulator was only producing 4.117 V. I was using a 5 V supply because I wanted to use only 1 power supply (I was using a two-channel power supply that had a fixed 5V output to produce the +15, -15, ground, and 5 V I need for my photosensor circuit box). After seeing this, I got a second power supply and am now using 10V to as an input for the regulator to produce 4.961V. I found that from a voltage range of 10V to 15 V, the regulator produced a steady 4.961 V supply. I have decided to use 10V as an input. My newly-grounded voltage regulator did not smoke or get hot at 10V.
After several more debugging trials (my LED was still not lighting up, according to the infared viewer), I learned another painful lesson. I learned DO NOT USE CLIP LEADS TO TEST CIRCUITS!!!! Initally, I was powering my circuit and making all of my connections between the photosensor head (2 photodiodes and 1 LED) with clip leads. This was a BAD IDEA BECAUSE CLIP LEADS ARE UNSTABLE AND IT IS VERY EASY TO SHORT A CIRCUIT IF THEY ACCIDENTALLY TOUCH! I did not realize this important lesson until my photosensor circuit was once again burning. Confused as to why my circuit was once again burning, I foolishly touched the voltage regulator. As you can see on the top voltage regulator in the photo below, my finger left its mark on the smoldering voltage regulator. As you cannot see the wincing on my face as I try to type this long elog, I will painfully type that the voltage regulator left its own mark on my finger (an ugly sore little welt). Suresh has taught me a valuable lesson: WHEN DEALING WITH SOMETHING OF QUESTIONABLE/UNKNOWN TEMPERATURE, USE YOUR NOSE, NOT YOUR FINGER TO DETERMINE IF THAT COMPONENT IS HOT!!!!

To make my circuit-testing safer, upon the suggestion of Suresh, I have since removed the clip leads and inserted a 12 pin IDC component (pictured below). There are 12 pins for the 6 inputs I will get from each of the 2 photosensor heads. I have requested orders for a 16 pin IDC connector, 15 pin Dsub male part, 15 pin Dsub feed-thru, 9 pin Dsub male part (2), and 9 pin Dsub feed-thru (2). After receiving these components, I should be able to safely test my circuit.

In the meanwhile, I can explore SimMechanics and try to figure out how to use the accelerometer |
4985
|
Mon Jul 18 21:06:32 2011 |
PSL Table Guardian | Omnistructure | PSL | Don't leave the PSL table open, unattended!!!!!!!!!!!!11111 |
I found the PSL table left open, and unattended again.
As far as I know, Jamie and Jenne (working on the LSC rack, so no lasers / optics work involved) have been the only ones in the IFO room for several hours now.
I'm going to start taking laser keys, or finding other suitable punishments. Like a day of lab cleanup chores or something. Seriously, don't leave the PSL table open if you're not actively working on it. |
4984
|
Mon Jul 18 20:59:19 2011 |
Jenne | Update | LSC | Big ol' mess |
[Jamie, Jenne]
We decided to take on the deceptively easy-sounding task of checking that the LSC whitening switching was happening as anticipated. We hoped to discover that when we clicked the "unwhitening" switches in FM1 of the LSC PDs, we would see the analog whitening turn on and off for the matching channel. That is what is supposed to happen.
Tragically, it is instead one big giant crazy disaster of a mess.
What we did:
Made a 24tapus (octopus like last time, except more...), with a 50kOhm resistor as our white noise source (instead of using a DAC channel and AWG).
We plugged our 24tapus into the 3 of 4 whitening boards on the LSC rack that are currently in use. One of the boards just has 8 terminators on the input, so we left that one alone for now.
We put the whitening gains to 0dB so that all the channels looked the same.
We looked at the PD _IN1 channels in DTT, and monitored which signals had whitening switching when we clicked the "unwhitening" buttons on the PD filter banks.
So far, we can find no rhyme or reason as to why some of the channels work (click unwhite on that PD, see that signal have whitening switching), and others don't. Some channels we just can't get to switch no matter what, others are just mis-mapped. There is no discernible pattern.
What we think (so far) is going on:
All of the cables from the PD demod boards are going to the Whitening board inputs, exactly as in Suresh's Diagram. The only difference is that Refl33, AS165 and Refl165 demod boards don't exist in the rack at this time.
The Whitening and AA boards in Suresh's Diagram labeled 0-7 are connected to Binary Output channels 0-7. This is a good thing.
The Whitening and AA boards in the diagram labeled 8-15 are connected to Binary Output channels 24-31. This is not so awesome.
This is all we are confident about at this time.
Next steps:
We are hoping that Ben has a secret stash (or can tell us who would) of LSC rack wiring diagrams. We would like to find out, without the pain of tracing wires and cables by hand, how the Binary I/O information gets through the cross-connect on the LSC rack up to the whitening boards.
We are leaving the 24tapus in place for now, so that we can carry on tomorrow, either with a wiring diagram in hand, or carefully tracing cables. |
4983
|
Mon Jul 18 15:51:04 2011 |
steve | Update | SUS | ETMY spectra of oplev error signal |
|
Attachment 1: ETMYopler.png
|
|
4982
|
Mon Jul 18 14:39:53 2011 |
Jenne | Update | Computers | LSC PD chans acquired |
There's too much tromping around, so I'm not going to actually measure PRC length right now, but I did set some channels to be acquired (POPDC, POXDC, POYDC) in addition to ASDC which was already acquired, so that I can look at the resonance fringes when I sweep the ABSL laser (hopefully later tonight....) |
4981
|
Mon Jul 18 10:24:48 2011 |
steve | HowTo | General | solvent bottles for cleaning optics |
Each bottle has matched seals. They are not interchangeable.
It is critical that the solvent do not reach the rubber bulb. Practice with the pipet.
In case of solvent touching the suction bulb: do not let the solvent go back into the bottle! Remove bulb, let it dry out and rinse pipet.
It is essential that the solvent bottle must be rinsed and refilled if it's content met with the rubber bulb.
Use glass syringe with SS needle in critical application: Hamilton ~0.1 ml |
Attachment 1: P1080090.JPG
|
|
4980
|
Sun Jul 17 18:23:23 2011 |
Jenne | Update | PSL | Found the PMC unlocked |
It was unlocked since ~4:30am. No idea why. It's relocked so I can try round N of measuring the PRC length. |
Attachment 1: PMCunlocked_17July2011.png
|
|
4979
|
Sat Jul 16 18:54:05 2011 |
Ishwita, Manuel | Configuration | Electronics | AA board |
We fixed the anti-aliasing board in its aluminum black box, the box couldn't be covered entirely because of the outgoing wires of the BNC connectors, so we drilled additional holes on the top cover to slide it backwards by 1cm and then screw it.
We had to fix the AA board box in rack 1X7, but there wasn't enough space, so we tried to move the blue chassis (ligo electro-optical fanout chassis 1X7) up with the help of a jack. We removed the blue chassis' screws but we couldn't move it up because of a piece of metal screwed above the blue chassis, then we weren't able to screw the two bottom screws again anymore because it had slided a bit down. Thus, the blue chassis (LIGO ELECTRO-OPTICAL FANOUT CHASSIS 1X7) is still not fixed properly and is sitting on the jack.
To accommodate the AA board (along with the panel-mounted BNC connectors) in rack 1X7 we removed the sliding tray (which was above the CPU) and fixed it there. Now the sliding tray is under the drill press.
|
Attachment 1: DSC_3236.JPG
|
|
Attachment 2: pic1.png
|
|
Attachment 3: DSC_3237.JPG
|
|
4978
|
Fri Jul 15 19:00:18 2011 |
dmass | Metaphysics | elog | Crashes |
Elog crashed a couple times, restarted it a couple times. |
4977
|
Fri Jul 15 17:42:21 2011 |
Sonali | Update | Green Locking | PSL layout for superposition of the PSL,ETMX and ETMY beams. |
The fibres carrying the beams from the ETMX as well as the ETMY have been routed to the PSL table now.
A part of the PSL beam has to be superposed on the fibre-outputs to obtain a beat signal. We have located a stray beam on the PSL(which is currently being dumped) which we plan to redirect for the same. The layout of the plan is attached herewith. |
Attachment 1: PSL_change1.pdf
|
|
Attachment 2: PSLnewspace1.pdf
|
|
4976
|
Fri Jul 15 16:14:00 2011 |
steve | Update | SUS | BS oplev error signal spectra |
|
Attachment 1: BS_opl_ersig.pdf
|
|
4975
|
Fri Jul 15 14:29:30 2011 |
Nicole | Summary | SUS | Photosensor Head Lessons |
You are right Jamie! Thank you for the correction! I will now use the Teflon sheet instead of the PCB piece.
The photodiodes do have three legs, but I imagined the third one lying on a different plane, since it is spaced apart from the two I have drawn.
I should include this third leg in my drawing? |
4974
|
Fri Jul 15 14:23:30 2011 |
Nicole | Summary | SUS | Photosensor Head Lessons |
Ah! I see! Thank you!
I should put the LEDs and photodiodes closer together so that more of the reflected light falls on the photodiodes and the photodiodes have a higher response.
Also the reflectivity of the mirror will be optimized if the incident light is normal to the mirror surface. We will be setting up the photosensor and mirror so that the LEDs
emit light normal to the mirror surface. During displacement, this light may be slightly off-normal but still close to normal incidence. We want the photodiodes to be close to the LED since we want
them to detect light that is close to the path of normal incidence (small angles of reflection). [Thanks to Jenne for helping me figure this one out!]
Thank you for the suggestion ^___^ |
4973
|
Fri Jul 15 13:48:56 2011 |
sonali | Update | Green Locking | Power of the AUX laser increased. |
Quote: |
Quote: |
The power of the beam which is being coupled into the optical fibre is measured to be between 159 mW to 164 mW (The power meter was showing fluctuating readings).
The power out of the beam coming out of the fibre far-end at the PSL table is measured to be 72 mW. Here, I have attached a picture of the beam paths of the ETMY table with the beams labelled with their respective powers.
|
For the phase locking or beat note measuring we only need ~1 mW. Its a bad idea to send so much power into the fiber because of SBS and safety. The power should be lowered until the output at the PSL is < 2 mW. In terms of SNR, there's no advantage to use such high powers.
|
Well,the plan is to put in a neutral density filter in the beam path before it enters the fibre. But before I could do that, I set up the camera on the PSL table to look at the fiber output . I will need it while I realign the beam after putting in the Neutral Density Filter. I have attached the ETMY layout with the Neutral Density filter in place herewith. |
Attachment 1: ETMY_after_fibre_coupling_labelled.pdf
|
|
4972
|
Fri Jul 15 09:25:02 2011 |
rana | Update | SUS | SUS oplev spectras |
In addition to the OL quadrants, you need to plot the OPLEV_PERROR and OPLEV_YERROR signals since these are the real signals we use for finding the mirror motion. If they're not in the Dataviewer, Jamie should add them as 256 Hz DAQ channels (using these names so that we have the continuity with the past). These DAQ channels correspond to the IN1 channels for the OL filter banks.
Also JPG are banned from the elog - you should put all of the plots into a single, multipage PDF file in honor of the new Wagonga. |
4971
|
Fri Jul 15 08:48:36 2011 |
Jamie | Summary | SUS | Photosensor Head Lessons |
Nicole: I thought we had decided to use teflon as the insulator between the PCB (yellow) and the LED/PDs? I don't think you should use another circuit board with copper on it. The copper will short the LED/PD heads to the metal box, which might be problematic.
Otherwise the design looks pretty good. I think the PDs have three leads each, yes? |
4970
|
Fri Jul 15 01:11:21 2011 |
Koji | Summary | SUS | Photosensor Head Lessons |
Rotate the PDs and the LED so that you can put them as close as possible.
This is to increase the sensitivity of the sensor. Think why the closer the better. |
4969
|
Thu Jul 14 20:24:32 2011 |
Nicole | Summary | SUS | Photosensor Head Lessons |
Today I tested the photosensor head combination (2 Hamamatsu S5971 photodiodes and 1 Hamamatsu L9337 LED). I discovered that I had burnt out the LED and the photodiodes when I soldered them to the PCB board.
After looking up soldering information on Hamamatsu photodiodes, I learned that I need to solder at least 2 mm away from the head. I checked the pins of my burnt-out photodiodes and I had soldered 1.5 mm away from the head. To prevent this problem from happening again, Suresh suggested that I clip a lead onto photodiode/LED pin while I solder on connections to help dissipate some of the heat.
Today I was able to get a single photodiode (not attached to the PCB) to measure light emitted from an LED and I observed how voltage fluctuated as I moved the photodiode around the LED.
Suresh and Jamie also helped me to fix my photosensor head design (to make it more electrically-stable). Originally, I had planned to solder the LED and photodiodes onto a PCB and to mount that PCB to the front of a small metal Pomona Electronics box (with a whole cut out for the photodiodes and LED) using spacers, screws, and nuts. However, the PCB I am using to solder on the LED and photodiodes has metal connections that may cause problems for the LED and photodiodes lying on the surface. Now, the plan is to have the LED and photodiodes mounted to the PCB with an insulatory PCB in between. Below is an explanatory picture. I will determine the placement of the LED and photodiodes after making screws holes in the two PCBs to attach to the metal face of the box. I want to attach the screw holes first to make sure that the PCBs (and attached photosensor) are centered.

|
4968
|
Thu Jul 14 17:34:35 2011 |
Ishwita, Manuel | HowTo | WienerFiltering | Wiener-Hopf equations |
Since we are using Wiener filtering in our project, we studied the derivation of Wiener-Hopf equations. Whatever we understood we have written it as a pdf document which is attached below... |
Attachment 1: derivwf.pdf
|
|
4967
|
Thu Jul 14 15:27:08 2011 |
steve, | Update | SUS | SUS oplev spectras |
Quote: |
Quote: |
I repeated the BS oplev spectrum today and I do not understand why it does look different. I did it as Kiwamu describes it in entry#4948 The oplev servo was left ON!
|
It is working today! Finally I repeated the BS spectra, that we did with Kiwamu last week
|
The optical levers were centered during these measurements without the reference of locked cavities. They have no reference value now.
SRM sus need some help. ITMX is showing pitch/yaw modes of the pendulum .....OSEM damping is weak? |
Attachment 1: BS_oplev.jpg
|
|
Attachment 2: PRM_oplev.jpg
|
|
Attachment 3: ITMX_oplev.jpg
|
|
Attachment 4: ETMX_oplev.jpg
|
|
Attachment 5: ETMY_oplev.jpg
|
|
Attachment 6: SRM_oplev.jpg
|
|
Attachment 7: ITMY_oplev_b.jpg
|
|
4966
|
Thu Jul 14 09:38:50 2011 |
steve, | Update | SUS | BS oplev spectra |
Quote: |
I repeated the BS oplev spectrum today and I do not understand why it does look different. I did it as Kiwamu describes it in entry#4948 The oplev servo was left ON!
|
It is working today! Finally I repeated the BS spectra, that we did with Kiwamu last week |
Attachment 1: BS_oplev.jpg
|
|
4965
|
Thu Jul 14 02:32:11 2011 |
sonali | Update | Green Locking | Power of the AUX laser increased. |
Quote: |
The power of the beam which is being coupled into the optical fibre is measured to be between 159 mW to 164 mW (The power meter was showing fluctuating readings).
The power out of the beam coming out of the fibre far-end at the PSL table is measured to be 72 mW. Here, I have attached a picture of the beam paths of the ETMY table with the beams labelled with their respective powers.
|
For the phase locking or beat note measuring we only need ~1 mW. Its a bad idea to send so much power into the fiber because of SBS and safety. The power should be lowered until the output at the PSL is < 2 mW. In terms of SNR, there's no advantage to use such high powers. |
4964
|
Wed Jul 13 12:24:46 2011 |
Nicole | Update | SUS | Weekly Update |
This week, I have been working on the photosensor circuit box. This photosensor box will contain the current-stabilizing power supply and
voltage readout for the two photosensors I plan to build.
Suresh helped to walk me through the design of the photosensor circuit (image below) so I now understand how the circuit works.
Jaimie helped me to reorganize the original circuit layout I had to make it easier to follow. I have now redone half of the circuit (enough for one LED and photodiode pair). I still need to put in the voltage-regulators to provide the + and - 15 V neeCto power the op-amps but I will do that after testing the circuit.
In order to test this preliminary circuit, I need to build the photosensor heads. Yesterday, Suresh helped me to open one of the professionally-build photosensors in the lab to understand how to arrange my photosensor heads. I now understand that I need to rigidly-mount the PCB to photosensor head box. I plan to use the PCB below. It will be sufficient for the lower-frequency range (below 10Hz) that I am interested in.

I would like to use a metal box like the one below to make each photosensor head. I looked in the lab last night for similar boxes but could not find one. Does anyone know where I can find a similar metal box?

I am now working on accelerometer. I am working on attaching these metal wires to the pins of the accelerometer so that I can use clip leads to power and extract voltage measurements from my circuit.
 |
4963
|
Tue Jul 12 17:30:24 2011 |
steve, | Update | SUS | BS oplev spectra |
I repeated the BS oplev spectrum today and I do not understand why it does look different. I did it as Kiwamu describes it in entry#4948 The oplev servo was left ON! |
Attachment 1: BS_oplev.jpg
|
|
4962
|
Tue Jul 12 11:52:54 2011 |
Jamie, Suresh | Update | LSC | LSC model updates |
The LSC model has been updated:
Binary outputs to control whitening filter switching
We now take the filter state bit from the first filter bank in all RF PD I/Q filter banks (AS55_I, REFL11_Q, etc) as the controls for the binary analog whitening switching on the RF PD I/Q inputs. The RF_PD part was also modified to output this control bit. The bits from the individual PDs are then combined into the various words that are written to the Contec BO part.
Channel mapping updated/fixed to reflect wiring specification
Yesterday Suresh posted an updated LSC wiring diagram, with correct channel assignments for the RF PD I/Q and DC inputs. Upon inspection of the physical hardware we found that some of LSC the wiring was incorrect, with I/Q channels swapped, and some of the PDs in the wrong channels. We went through and fixed the physical wiring to reflect the diagram. This almost certainly will affect the EPICS settings for some of the input channels, such as offsets and RD rotations. We should therefore go through all of the RF inputs and make sure everything is kosher.
I also fixed all of the wiring in the LSC model to also reflect the diagram.
Once this was all done, I rebuilt and restarted the LSC model, and confirmed that the anti-whitening filter banks in the PD input filter modules were indeed switching the correct bits. I'll next put together a script to confirm that the LSC PD whitening is switching as it should.
|
4961
|
Tue Jul 12 10:18:05 2011 |
Jamie | Update | CDS | C1:DAQ-FB0_C1???_STATUS indicators red, restored after controller restarts |
Yesterday I found the C1:DAQ-FB0_C1???_STATUS lights to be red for the SUS, MCS, SCX, and SCY controllers. I know this has something to do with model communication with the framebuilder, but I unfortunately don't remember exactly what it is. I decided to try restarting the affected models to see if that cleared up the problem. It did. After restarting c1scx, c1scy, c1sus, and c1mcs everything came back up green.
We need some better documentation about what all of these status indicators mean. |
4960
|
Mon Jul 11 14:03:37 2011 |
steve | HowTo | lore | how to visit your old lab |
Alberto is visiting us from Australia. He brought some terrific presents. It is going to be very demanding task to wait for the rest of the 40m team
to return from Wales to taste coffee: PNG Peaberry of Wagonga, Monsooned Malabar of Jindebah and Signature Blue Blend of Cosmorex. |
Attachment 1: P1080069.JPG
|
|
4959
|
Mon Jul 11 10:10:31 2011 |
Ishwita | Configuration | | AA board |
The AA board shown in attachment 1 will be used in the seismometer hardware setup. A cartoon of this setup is shown in attachment 2.
BNC connectors are required for the seismometer breakout boxes. So the four-pin LEMO connectors present in the AA board were removed and panel mount BNC connectors were soldered to it. Red and blue colored wires were used to connect the BNC connectors to the board. Red wire connects the center of the BNC connector to a point on the board and that connection leads to the third leg (+IN) of the IC U### and the blue wire connects the shield of the BNC connector to the second leg (-IN) of the IC U###.
All the connections (including BNC to the AA board and in the AA board to all the filters) were tested using a multimeter by the beeping method and it was found that channel 10 (marked as C10) had a wrong connection from the point where the red wire (+ve) was connected to the third leg (+IN) of IC U91 and channel 32 (marked as C32) had opposite connections meaning the blue wire is connected to the third leg (+IN) of IC U311 and red wire is connected to the second leg (-IN) of IC U311. |
Attachment 1: P7080305.JPG
|
|
Attachment 2: seismometers.png
|
|
4958
|
Fri Jul 8 20:50:49 2011 |
sonali | Update | Green Locking | Power of the AUX laser increased. |
The ETMY laser was operating at 1.5 A current and 197 mW power.
For the efficient frequency doubling of the AUX laser beam at the ETMY table, a higher power is required.
Steve and I changed the current level of the laser from 1.5 A to 2.1 A in steps of 0.1 A and noted the corresponding power output . The graph is attached here.
The laser has been set to current 1.8 Amperes. At this current, the power of the output beam just near the laser output is measured to be 390 mW.
The power of the beam which is being coupled into the optical fibre is measured to be between 159 mW to 164 mW (The power meter was showing fluctuating readings).
The power out of the beam coming out of the fibre far-end at the PSL table is measured to be 72 mW. Here, I have attached a picture of the beam paths of the ETMY table with the beams labelled with their respective powers.
Next we are going to adjust the green alignment on the ETMY and then measure the power of the beam.
At the output end of the fibre on the PSL, a power meter has been put to dump the beam for now as well as to help with the alignment at the ETMY table. |
Attachment 1: Graph3.png
|
|
Attachment 2: ETMY_beam_powers.png
|
|
4957
|
Fri Jul 8 19:50:19 2011 |
Suresh | Update | RF System | LSC rack channel assignment |
[Jamie, Suresh]
We looked at the ADC channel assignments in the LSC model and wanted to make sure that the LSC rack wiring and the LSC model are in agreement with each other. So the plan is to wire the rack as shown below. I will also post this file on svn so that we can keep it updated in case there are changes.

|
4956
|
Fri Jul 8 09:53:49 2011 |
Nicole | Summary | SUS | Summer Progress Report 1 |
A copy of my summer progress report 1 has been uploaded to ligodcc 7/711 and I have just added a copy to the TTsuspension wiki
PDF copy of Summer Progress Report |
4955
|
Thu Jul 7 15:34:44 2011 |
Jenne | Summary | LSC | LSC Whitening Filters have been fit |
Quote: |
During the daytime either tomorrow or Friday I'll adjust the actual dewhitening filters to match the measured zpk values.
|
I made a handy-dandy table showing the zpk values for each whitening filter in the wiki: New whitening filter page
Next on the whitening filter to-do list: actually put these values into the dewhitening filters in foton. |
4954
|
Thu Jul 7 14:27:16 2011 |
Suresh | Update | IOO | something wrong ? : Power incident on REFL11 and REFL55 |
Just tying up a loose end. The next day Kiwamu and I checked to see what the trouble was. We concluded that the PRM had not moved during my measurement though I had 'Misaligned' it from the medm screen. So all the power levels measured here were with the PRM aligned. The power level change was subsequently measured and e-logged
Quote: |
The measured change in the REFL DC power with and without PRM aligned seems unacceptably small. Something wrong ?
The difference in the power with and without PRM aligned should be more than a factor of 300.
[difference in power] = [single bounce from PRM] / [two times of transmission through PRM ]
= (1-T) / T^2 ~ 310,
where T is the transmissivity of PRM and T = 5.5% is assumed in the calculation.
Also the reflectivity of MICH is assumed to be 1 for simplicity.
Quote from #4894 |
We now have (with the PRM misaligned):
REFL11: Power incident = 7.60 mW ; DC out = 0.330 V => efficiency = 0.87 A/W
REFL55: Power incident = 23 mW ; DC out = 0.850 V => efficiency = 0.74 A/W
and with the PRM aligned::
REFL11: DC out = 0.35 V => 8 mW is incident
REFL55: DC out = 0.975 V => 26 mW is incident
|
|
|
4953
|
Thu Jul 7 13:57:28 2011 |
steve | HowTo | SAFETY | how not to |
I found baked allen keys on the top of the clean optics cabinet. Somewhat heavy box that can come down in an earthquake on our heads.
NOTHING SHOULD GO ON THE TOP OF THE CABINETS OR RACKS except small plastic boxes that storing our clean clothing. |
4952
|
Thu Jul 7 10:25:34 2011 |
steve | HowTo | General | cleaning out refregerator |
Please ask the owner unless it is rotten. Do not put food into garbage can inside. Take them outside so you are not inviting ants ! |
4951
|
Thu Jul 7 02:23:59 2011 |
Jenne | Summary | LSC | LSC Whitening Filters have been fit |
I have fit all of the LSC whitening filters using vectfit4.m
All the data is in my folder ..../users/jenne/LSC_WhiteningTest_29June2011/
The zpk info is saved with each plot of the fit. The pdfs are kind of huge to stitch together (or rather my computer doesn't want to do it), so I'll just post a representative one for now.

During the daytime either tomorrow or Friday I'll adjust the actual dewhitening filters to match the measured zpk values. |
4950
|
Wed Jul 6 23:53:38 2011 |
kiwamu | Summary | General | minutes of 40m meeting |
|
Attachment 1: meeting.pdf
|
|
4949
|
Wed Jul 6 23:03:57 2011 |
kiwamu | Update | LSC | modified locking scripts |
[Jenne / Kiwamu]
Last night we modified the locking scripts, that were called from C1IFO_CONFIGURE.adl, to adapt them to the new "PRCL" and "SRCL" convention.
So far they work fine and quitted dumping some error messages about inexistence of these channel names.
P.S. The locking scripts have been summarized on the 40m wiki
Quote from #4912 |
- Now the power and signal recycling cavity lengths are named "PRCL" and "SRCL" in stead of three letter names without "L".
We should change the locking script to accomodate these changes.
|
|
4948
|
Wed Jul 6 18:15:26 2011 |
kiwamu | Update | SUS | Re: BS oplev spectra |
[Steve / Kiwamu]
Motivation:
Since the oplevs were the ones we haven't carefully tested, so the oplevs need to be checked.
This checking is also a part of the suspension optimizations (see the minutes of the last 40m meeting).
In this work Steve will check two things for all the oplevs :
1. Noise level including the dark noise, electrical noise and ADC noise to just make sure that the noise are blow the signal levels below ~ 30Hz.
2. The spectra of the signals to make sure there are no funny oscillations and unexpected structures
Measurement :
To check the things listed above, we take two kinds of oplves' spectra :
1. "dark noise" when the He-Ne beam is blocked.
2. "signals" when the optics are damped by only OSEMs
We did these checks on the BS oplev today (see the last entry).
All of them are fine, for example the dark noise (including electrical noise and ADC noise) are below the signal levels.
And no oscillation peak was found. Steve will go through all of the oplevs in this way.
Quote from #4947 |
Healthy BS oplev
|
|
4947
|
Wed Jul 6 16:44:37 2011 |
steve, kiwamu | Update | SUS | BS oplev spectra |
Healthy BS oplev |
Attachment 1: BS.jpg
|
|
4946
|
Wed Jul 6 15:32:32 2011 |
Jamie | Update | SUS | Re : More normalization of all sus controllers |
So after talking to Kiwamu about it, I understand now that since the damping loops need all of this extra gain when the high-pass corner is moved up, it's more convenient to put that gain in the control filter itself, rather than having to crank the overall DC gain up to some inconveniently high value. |
4945
|
Wed Jul 6 11:45:20 2011 |
Jamie | Update | SUS | More normalization of all sus controllers |
Quote |
I'm attaching a screenshot of some of the problems I see so far with MC3.
|
I tried to fix all of the problems that I could identify in this screen shot:
- Fixed the TO_COIL output filter matrix screen to correctly point to the matrix element filter screens (all SUS)
- Removed MCL sections from SUS_XXX_POSITION screens, except for MC2. I also modified the _POSITION screens for the ETMs to refer to ALS instead of MCL.
- Zeroed out all of the lockin gains in the TO_COIL matrices (MC SUS)
- Made sure all whitening filter were ON (all SUS)
- Made sure all cts2um calibration filters were on (all SUS)
- Made sure all oplev servos were on (all SUS)
|
4944
|
Wed Jul 6 10:35:35 2011 |
Jamie | Update | SUS | Re : More normalization of all sus controllers |
Quote: |
We found the 30 Hz high pass filters had lower gain than what they used to be at low frequcnies.
So we increased the gain of the high pass filters called '30:0.0' by a factor of 10 to have the same gain as before.
|
I'm not convinced that this is what you want to do, or at least I wouldn't do it this way. The "k" in the zpk filter was set such that the filter had unity gain above the high-pass cut-off frequency. For a 30 Hz high-pass the k needs to be a factor of 10 smaller than it would be for a 3 Hz high-pass to achieve this high frequency unity gain.
As it is now these HP filters have 20 dB of gain above 30 Hz. If the open loop transfer function needs to more gain I would have done that by adjusting the overall DC gain of the filter bank, not by increasing the gain in this one filter. Maybe you guys have been doing it differently, though. Or maybe I'm just completely off base. |
4943
|
Wed Jul 6 02:12:36 2011 |
kiwamu | Update | SUS | Re : More normalization of all sus controllers |
[Jenne / Rana/ Kiwamu]
We found the 30 Hz high pass filters had lower gain than what they used to be at low frequcnies.
So we increased the gain of the high pass filters called '30:0.0' by a factor of 10 to have the same gain as before.
Now all the suspension shows some kind of damping. Needs more optimizations, for example Q-adjustments for all the suspensions...
Quote from #4942 |
This is getting closer, but with the whitening left OFF and the cts2um filter also OFF, none of the suspensions are working correctly.
|
|
4942
|
Tue Jul 5 21:26:51 2011 |
rana | Update | SUS | More normalization of all sus controllers |
This is getting closer, but with the whitening left OFF and the cts2um filter also OFF, none of the suspensions are working correctly. I'm shutting down all the watchdogs until someone gets around to setting the damping gains and filters correctly.
I'm attaching a screenshot of some of the problems I see so far with MC3.
I'm going to try to get the MC suspensions working OK for tonight so that we can use them for the PRMI locking work.
Update #1: None of the MC SUS DAQ channels are found by dataviewer....SUS debugging speed reduced by 10x. Tue Jul 05 21:38:17 2011
Update #2: POS/PIT/YAW BIAS sliders now seem to work, but are ~1000x too weak to do anything. Tue Jul 05 21:41:38 2011
|
Attachment 1: Screenshot-1.png
|
|